OpenStudy (anonymous):

x+2 over 2 - 4x-12 over 5= 1 ???

OpenStudy (anonymous):

are you trying to solve for x?

OpenStudy (anonymous):

x=-52/21

OpenStudy (anonymous):

thank you but can you please explain how you solved it as well ??

OpenStudy (campbell_st):

multiply every term by 10... this will eliminate the fractions... \[\frac{10(x + 2)}{2} - \frac{10(4 -12)}{5} = 10 \times 1\] simplify, distribute, collect like terms then solve

OpenStudy (campbell_st):

oops should be 4x not 4

OpenStudy (anonymous):

ill do that thank you !

OpenStudy (anonymous):

the answer is x=8 right ?

OpenStudy (campbell_st):

ummm substituting (8 + 2)/2 - (4(8) -12)/5 = 5 - 4 looks good